OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). y = 2x – 7 y = −x − 2 y = 2x – 2 y = −x − 7

OpenStudy (anonymous):

So the first step for this problem is recognizing that the slope of the original line is 2 (remember that y = mx + b where "m" is the slope). Next, remember that two lines that are parallel share the same slope. Because of this, options 2 and 4 can be crossed off, as they do not have a slope of 2. Next, check which of the last 2 possible answers contains the point (5,3) by plugging in x=5 and y=3. One of them does not work!

OpenStudy (anonymous):

the second and last one is supposed to say y=-1/2x-2 and the last one is y=-1/2x-7

OpenStudy (anonymous):

does that make a difference ?

OpenStudy (mathmate):

Not if the line is supposed to be parallel to y = 2x + 2. The modified lines are perpendicular to y = 2x + 2, so are rejected as well.

OpenStudy (anonymous):

No it doesn't because those don't have the same slope (2), so you can cross those off. It actually makes the problem a lot easier, as the only equation that has the same slope as the original line is the first possible one!

OpenStudy (anonymous):

so is the answer A?

OpenStudy (mathmate):

Can you explain how you got A?

OpenStudy (anonymous):

Plug in (5,3) into answer A and see if it satisfies the equation. If it does (since the slope is the same as well), you're right!